OpenStudy (anonymous):

When the equation 3x - 4y = 24 is written in slope-intercept form, what is the value of m?

OpenStudy (anonymous):

slope intercept form is y = mx + b. 3x - 4y = 24 -4y = -3x + 24 (multiply by -1) 4y = 3x - 24 (divide by 4) y = 3/4x - 6 m = 3/4

OpenStudy (anonymous):

Thank You For Explaining , I understand clearly now

OpenStudy (anonymous):

so what about Solve V equals one third Bh for B.

OpenStudy (anonymous):

YW :)

OpenStudy (anonymous):

V=1/3Bh solve for B

OpenStudy (anonymous):

I am very much confused

OpenStudy (anonymous):

V = 1/3bh for B multiply by 3 to get rid of the fraction 3V = bh divide by h 3V/h = b ....I think or... V = 1/3bh divide by 1/3h V/(1/3)h = b not sure which one to use

OpenStudy (andriod09):

Lele, to solve for B using \[v=\frac {1}{3}Bh\] then you would do this: \[\frac {1}{3}B=vh\]